From the Contingency Table below, showing the applicants to a Summer Intern program, P(LA & MS) = :

UNDERGRAD

GRADUATE

TOTAL

MISSISSIPPI

40

20

60

ALABAMA

20

50

70

LOUISIANA

40

30

70

TOTAL

100

100

200

(Your answer should include a % sign without decimal points, e.g. 50%)
Answer #1

Answer

we will use the formula

P(A and B) = P(A) * P(B) for independent events

LA (LOUISIANA) and MS(MISSISSIPPI) are independent events because they have nothing in common.

So, P(LA & MS) = P(LA)*P(MS)

where P(MS) = (number of students in MS)/(total) = 60/200

and P(LA) = (number of students in LA)/(total) = 70/200

this implies

P(LA & MS) = (60/200)*(70/200)

= 4200/40000

= 0.105

converting to percent

= 0.105*100

= 10.5

= 11%